# PrismAudio: Decomposed Chain-of-Thoughts and Multi-dimensional Rewards for Video-to-Audio Generation

### v2

@@ 1 Introduction
Recent V2A advances (Zhang et al., 2024; Xing et al., 2024; Wang et al., 2024b) have evolved from direct synthesis to increasingly rich conditioning mechanisms. Early approaches like V2A-Mapper (Wang et al., 2024a) and Diff-Foley (Luo et al., 2023) rely solely on visual inputs, using embedding projection and contrastive alignment, respectively, but suffer from limited semantic precision and controllability. Subsequent methods (Chen et al., 2025; Mo et al., 2024; Tian et al., 2025) incorporate explicit text conditioning—MovieGen Audio (Polyak et al., 2024) via cross-attention in diffusion transformers, MMAudio (Cheng et al., 2024a) through multimodal transformers—yet remain opaque “black boxes” despite improved control. Most recently, ThinkSound (Liu et al., 2025b) pioneers Chain-of-Thought (CoT) reasoning (Wei et al., 2022) using multimodal LLMs (MLLMs) (Achiam et al., 2023; Cheng et al., 2024b; Chu et al., 2024), decomposing V2A into structured planning followed by audio rendering. This explicit reasoning significantly enhances interpretability and narrative coherence by making the generation process transparent and controllable.
@@
PrismAudio builds upon a CoT-aware audio foundation model employing a Multimodal Diffusion Transformer backbone with flow matching. Applying RL to diffusion models poses computational challenges (Xue et al., 2025; Li et al., 2025). While Group Relative Policy Optimization (GRPO) (Shao et al., 2024) shows promise for human preference alignment, current implementations like Flow-GRPO (Liu et al., 2025c) require Stochastic Differential Equation (SDE) sampling at every denoising step, creating substantial training overhead due to full-step sampling requirements for policy ratio computation. We propose Fast-GRPO, employing a hybrid ODE-SDE strategy—applying SDE sampling only to a subset of steps for stochastic exploration while using deterministic Ordinary Differential Equation (ODE) sampling elsewhere. Fast-GRPO enables efficient multi-dimensional CoT-RL optimization without compromising generation quality.
Evaluating practical V2A capabilities demands a rigorous benchmark covering realistically diverse and challenging scenarios; yet existing V2A benchmarks such as VGGSound (Chen et al., 2020) and Kling-Audio-Eval (Wang et al., 2025) fail to meet the requirements (see Appendix C.2 for detailed analysis). We therefore introduce AudioCanvas, featuring: (1) high modality alignment through rigorous off-screen sound filtering, (2) advanced scene complexity with 300 single-event classes and 501 multi-event samples across diverse scenes, and (3) precise audio captions with rich, structured CoT reasoning enabling comprehensive evaluation of semantic consistency, temporal synchrony, aesthetic quality, and spatial accuracy. Our main contributions are as follows:
@@ 2 Related Work
CoT Reasoning for Audio Generation. Large Language Models (LLMs) (Guo et al., 2025; Team et al., 2024b; Yang et al., 2025) have demonstrated remarkable reasoning capabilities through CoT prompting (Wei et al., 2022), enabling complex problem decomposition via intermediate reasoning steps. This paradigm has been extended to MLLMs, which integrate visual and audio understanding with linguistic reasoning (Achiam et al., 2023; Lin et al., 2023; Alayrac et al., 2022). The related works on V2A generation are summarized in Appendix A. Early V2A approach (Xie et al., 2024) uses vision-language models for video captioning, then employs text-to-audio models for synthesis. Recent works adopt video-audio-language MLLMs like VideoLLaMA2 (Cheng et al., 2024b) for structured CoT planning. ThinkSound (Liu et al., 2025b) exemplifies this by generating detailed audio descriptions before synthesis, improving semantic consistency and narrative coherence. However, existing MLLM-based approaches employ monolithic planning that cannot handle competing objectives or provide targeted optimization for distinct perceptual dimensions. Our work decomposes monolithic planning into four specialized CoT modules—Semantic, Temporal, Aesthetic, and Spatial—each providing focused reasoning with corresponding reward signals for multi-dimensional preference optimization.
Reinforcement Learning for Diffusion Models. Reinforcement Learning has achieved remarkable success in LLMs through RLHF (Ouyang et al., 2022; Bai et al., 2022), demonstrating the crucial role of aligning model outputs with human preference beyond likelihood maximization. Recent works have explored RL applications to diffusion models for preference alignment. Early approaches (Fan & Lee, 2023; Black et al., 2023; Fan et al., 2023) optimize diffusion score functions through policy gradient methods, while  Wallace et al. (2024) introduces DPO (Rafailov et al., 2023) to diffusion models for direct learning from human feedback. Most recently, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) based approaches have advanced RL-enhanced diffusion models. Flow-GRPO (Liu et al., 2025c) and DanceGRPO (Xue et al., 2025) introduce GRPO to flow matching models (Lipman et al., 2022), enabling divergent sampling by transforming ODEs into equivalent SDEs with reduced variance through group-based optimization. However, existing RL approaches for generation primarily focus on single-objective optimization and have not been extended to V2A generation, which expects multi-dimensional alignment across semantic, temporal, aesthetic, and spatial aspects. Our work pioneers the application of flow-matching GRPO to V2A generation with specialized multi-dimensional reward decomposition.
@@ 3.1 CoT-Aware Audio Foundation Model
We build our audio foundation model on the diffusion transformer backbone (Peebles & Xie, 2023; Liu et al., 2023b) with flow matching that takes video inputs and text conditioning to generate audio outputs. It undergoes standard pre-training on large-scale video-audio pairs to establish basic generation capabilities. While this architecture provides a solid foundation for V2A generation, it has two critical limitations that hinder effective multi-dimensional CoT reasoning: insufficient video understanding for complex and diverse scenarios, and limited text processing capabilities for structured reasoning content. Therefore, we enhance the ThinkSound architecture (Liu et al., 2025b) with the following two modifications to facilitate multi-dimensional CoT reasoning.
VideoPrism for Enhanced Video Understanding. Most existing V2A models, including ThinkSound, adopt CLIP-based image encoders (Radford et al., 2021) that process video frames independently as static images. This approach lacks comprehensive video understanding and fails to handle complex, diverse video scenarios in real-world applications. We replace CLIP with VideoPrism (Zhao et al., 2024), a state-of-the-art (SOTA) video encoder pre-trained on large-scale video data. VideoPrism employs a unified vision transformer architecture specially designed for video understanding, capturing rich semantic representations of objects, actions, and environmental contexts that are crucial for our multi-dimensional reasoning modules.
T5-Gemma for CoT-Aware Text Encoding. Our CoT modules produce analytical text containing logical structures, causal relationships, and multi-faceted reasoning patterns that require sophisticated language understanding; yet the standard T5 encoders in ThinkSound struggle with the complex, structured reasoning text generated by our CoT modules. Hence, we make another essential enhancement by upgrading the T5 encoder to T5-Gemma (Zhang et al., 2025). T5-Gemma adapts the reasoning capabilities of decoder-only LLMs into an efficient encoder-decoder architecture, effectively enabling proper conditioning on our structured CoTs for the generation model with its stronger reasoning comprehension capabilities.
@@ 3.2 Decomposing Multi-Dimensional CoT Reasoning
To achieve superior reasoning capabilities simultaneously across all four dimensions, we first employ Gemini 2.5 Pro (Comanici et al., 2025) for CoT data construction, leveraging its outstanding multimodal understanding and strong reasoning capabilities. Next, using this high-quality training data, we fine-tune the highly competitive open-source video language model VideoLLaMA2 (Cheng et al., 2024b) to generate four specialized CoTs: Semantic CoT identifies audio events and their characteristics from audio-visual content; Temporal CoT determines the sequential ordering of audio events; Aesthetic CoT focuses on audio quality aspects like naturalness and fidelity; and Spatial CoT analyzes sound positioning, including directional placement and distance for proper spatialization. The four specialized CoTs are then concatenated in this order to form the multi-dimensional CoT (as depicted in Figure 1) and used as enhanced, structured text conditioning to fine-tune our audio foundation model (Section 3.1), enabling the model to learn from explicit reasoning patterns and acquire better generalization by understanding the underlying logic behind audio-visual correspondences.
@@ 3.3.1 Multi-dimensional Reward Functions
To address this limitation, we design four specialized reward functions that align with our CoT dimensions: Semantic Reward, measured by MS-CLAP (Elizalde et al., 2024), a commonly used audio-text alignment model for evaluating content similarity; Temporal Reward, assessed via Synchformer (Iashin et al., 2024), a highly competitive model specifically designed to detect audio-visual synchrony; Aesthetic Reward, which uses the leading-edge assessor for audio aesthetic quality, Meta Audiobox Aesthetics (Tjandra et al., 2025), as a no-reference model trained to predict human Mean Opinion Scores (MOS); and Spatial Reward, which employs the high-performing StereoCRW (Chen et al., 2022) to verify directional positioning accuracy. This multi-objective approach allows for a balanced and comprehensive optimization across all key perceptual dimensions.
@@ 3.3.2 Fast-GRPO with Random-Window Exploration
To align the audio foundation model with the multi-dimensional human preference, we adopt GRPO for its stability. While generation of our flow matching model is inherently deterministic (an ODE), it can be equivalently formulated as a stochastic process (an SDE) that enables RL-based optimization (see Appendix B.1 for details). Prior works (Xue et al., 2025; Liu et al., 2025c) construct the Markov Decision Process (MDP) within the GRPO training process by applying this SDE formulation across the entire denoising trajectory. This “pure SDE” approach, however, forces GRPO to evaluate the policy at every step, creating a significant efficiency bottleneck.
@@
where $A^{i}$ is the group-normalized advantage for the $i$-th sample (Eq. 6). This design reduces the policy-model NFE (Number of Function Evaluations) from $T$ to $w$ per sample, yielding a near-linear complexity of GRPO training. We notice that some contemporaneous works, such as MixGRPO (Li et al., 2025), also propose hybrid ODE-SDE. Considering the concurrency of research and their differences from Fast-GRPO on window design, modalities, and scopes, our Fast-GRPO is a valid innovation for enabling efficient multi-dimensional RL training of diffusion models.
@@ 4.1 Experimental Setup
Evaluation Metrics. We conduct comprehensive evaluations using both objective and subjective metrics to assess the four key perceptual dimensions. For objective evaluation, we adopt established metrics across multiple dimensions. Following ThinkSound, we employ CLAP score for text-audio semantic alignment, DeSync measured by Synchformer for video-audio temporal synchrony, Fréchet Distance (FD(Kilgour et al., 2018) in the VGGish feature space, and Kullback-Leibler (KL) Divergence (Copet et al., 2024) based on predictions from the PaSST model for audio distribution similarity Liu et al. (2025b). For spatial accuracy of generated stereo audio, we adopt GCC MSE and CRW MSE Sun et al. (2024) to evaluate both difference of arrival (DoA) and interaural time difference (ITD). To measure aesthetic quality, we evaluate production quality (PQ), production complexity (PC), content enjoyment (CE), and content usefulness (CU) scores from Audiobox-Aesthetics (Tjandra et al., 2025). For subjective evaluation, we employ Mean Opinion Score (MOS) across two complementary dimensions: MOS-Q (Quality) evaluates the aesthetic quality and audio fidelity of generated audio, while MOS-C (Consistency) evaluates the comprehensive alignment between audio and video, encompassing semantic consistency, temporal synchrony, and spatial accuracy. More details of evaluation metrics are in Appendix F.
Implementation Details are in Appendix D. Since the multi-dimensional CoT fine-tuning and the RL post-training are based on VGGSound, evaluations on the VGGSound test set are in-domain evaluations. Competitive baselines include Frieren (16k, mono) (Wang et al., 2024b), V2A-Mapper (16k, mono) (Wang et al., 2024a), AudioX (44k, stereo) (Tian et al., 2025), HunyuanVideo-Foley (44k, mono) (Shan et al., 2025), MMAudio (44k, mono) (Cheng et al., 2024a), and ThinkSound (44k, stereo) (Liu et al., 2025b).
@@ Appendix A Related Work on Video-to-Audio Generation
V2A generation has recently gained significant attention with advances in multimodal AI systems (Siméoni et al., 2025; Brooks et al., 2024; Team et al., 2024a). Current V2A methods predominantly employ latent diffusion models (Xing et al., 2024; Liu et al., 2025a; Wang et al., 2024b), while some explore autoregressive token-based approaches. Diff-Foley (Luo et al., 2023) and VTA-LDM (Xu et al., 2024) represent standard latent diffusion models conditioned on video features, while FoleyGen (Mei et al., 2024b) and V-AURA (Viertola et al., 2025) frame conditional audio generation as next-token prediction using visual features. Early methods focus on improving semantic consistency through better video representations. Works like CAVP (Luo et al., 2023) and CLIP4CLIP (Luo et al., 2022) employ contrastive learning for video encoding, while adapter-based approaches like FoleyCrafter (Zhang et al., 2024) build upon pre-trained text-to-audio models (Liu et al., 2023a, 2024b, 2024a) for enhanced controllability. Recent advances introduce explicit multimodal conditioning to address semantic limitations. MovieGen Audio (Polyak et al., 2024) conditions on both text and video to generate video-aligned audio, achieving substantial progress and inspiring subsequent video-text-audio generation works (Cheng et al., 2024a; Chen et al., 2025; Shan et al., 2025; Tian et al., 2025). Most recently, ThinkSound (Liu et al., 2025b) innovatively introduces CoT reasoning via MLLMs, replacing simple text prompts with structured reasoning that significantly improves interpretability and narrative coherence. However, existing V2A methods suffer from objective entanglement—optimizing competing perceptual goals through a single reconstruction loss—and lack human preference alignment beyond textual matching. In contrast, we propose PrismAudio, the first reinforcement learning framework for V2A generation with specialized multi-dimensional CoT-reward correspondence to address these fundamental limitations.

### v3

@@ 1 Introduction
Recent V2A advances (zhang2024foleycrafter; xing2024seeing; wang2024frieren) have evolved from direct synthesis to increasingly rich conditioning mechanisms. Early approaches like V2A-Mapper (wang2024v2a) and Diff-Foley (luo2023diff) rely solely on visual inputs, using embedding projection and contrastive alignment, respectively, but suffer from limited semantic precision and controllability. Subsequent methods (chen2025video; mo2024text; tian2025audiox) incorporate explicit text conditioning—MovieGen Audio (polyak2024movie) via cross-attention in diffusion transformers, MMAudio (cheng2024taming) through multimodal transformers—yet remain opaque “black boxes” despite improved control. Most recently, ThinkSound (liu2025thinksound) pioneers Chain-of-Thought (CoT) reasoning (wei2022chain) using multimodal LLMs (MLLMs) (achiam2023gpt; cheng2024videollama; chu2024qwen2), decomposing V2A into structured planning followed by audio rendering. This explicit reasoning significantly enhances interpretability and narrative coherence by making the generation process transparent and controllable.
@@
PrismAudio builds upon a CoT-aware audio foundation model employing a Multimodal Diffusion Transformer backbone with flow matching. Applying RL to diffusion models poses computational challenges (xue2025dancegrpo; li2025mixgrpo). While Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) shows promise for human preference alignment, current implementations like Flow-GRPO (liu2025flow) require Stochastic Differential Equation (SDE) sampling at every denoising step, creating substantial training overhead due to full-step sampling requirements for policy ratio computation. We propose Fast-GRPO, employing a hybrid ODE-SDE strategy—applying SDE sampling only to a subset of steps for stochastic exploration while using deterministic Ordinary Differential Equation (ODE) sampling elsewhere. Fast-GRPO enables efficient multi-dimensional CoT-RL optimization without compromising generation quality.
Evaluating practical V2A capabilities demands a rigorous benchmark covering realistically diverse and challenging scenarios; yet existing V2A benchmarks such as VGGSound (chen2020vggsound) and Kling-Audio-Eval (wang2025kling) fail to meet the requirements (see Appendix C.2 for detailed analysis). We therefore introduce AudioCanvas, featuring: (1) high modality alignment through rigorous off-screen sound filtering, (2) advanced scene complexity with 300 single-event classes and 501 multi-event samples across diverse scenes, and (3) precise audio captions with rich, structured CoT reasoning enabling comprehensive evaluation of semantic consistency, temporal synchrony, aesthetic quality, and spatial accuracy. Our main contributions are as follows:
@@ 2 Related Work
CoT Reasoning for Audio Generation. Large Language Models (LLMs) (guo2025deepseek; team2024gemma; yang2025qwen3) have demonstrated remarkable reasoning capabilities through CoT prompting (wei2022chain), enabling complex problem decomposition via intermediate reasoning steps. This paradigm has been extended to MLLMs, which integrate visual and audio understanding with linguistic reasoning (achiam2023gpt; lin2023video; alayrac2022flamingo). The related works on V2A generation are summarized in Appendix A. Early V2A approach (xie2024sonicvisionlm) uses vision-language models for video captioning, then employs text-to-audio models for synthesis. Recent works adopt video-audio-language MLLMs like VideoLLaMA2 (cheng2024videollama) for structured CoT planning. ThinkSound (liu2025thinksound) exemplifies this by generating detailed audio descriptions before synthesis, improving semantic consistency and narrative coherence. However, existing MLLM-based approaches employ monolithic planning that cannot handle competing objectives or provide targeted optimization for distinct perceptual dimensions. Our work decomposes monolithic planning into four specialized CoT modules—Semantic, Temporal, Aesthetic, and Spatial—each providing focused reasoning with corresponding reward signals for multi-dimensional preference optimization.
Reinforcement Learning for Diffusion Models. Reinforcement Learning has achieved remarkable success in LLMs through RLHF (ouyang2022training; bai2022constitutional), demonstrating the crucial role of aligning model outputs with human preference beyond likelihood maximization. Recent works have explored RL applications to diffusion models for preference alignment. Early approaches (fan2023optimizing; black2023training; fan2023dpok) optimize diffusion score functions through policy gradient methods, while  wallace2024diffusion introduces DPO (rafailov2023direct) to diffusion models for direct learning from human feedback. Most recently, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) based approaches have advanced RL-enhanced diffusion models. Flow-GRPO (liu2025flow) and DanceGRPO (xue2025dancegrpo) introduce GRPO to flow matching models (lipman2022flow), enabling divergent sampling by transforming ODEs into equivalent SDEs with reduced variance through group-based optimization. However, existing RL approaches for generation primarily focus on single-objective optimization and have not been extended to V2A generation, which expects multi-dimensional alignment across semantic, temporal, aesthetic, and spatial aspects. Our work pioneers the application of flow-matching GRPO to V2A generation with specialized multi-dimensional reward decomposition.
@@ 3.1 CoT-Aware Audio Foundation Model
We build our audio foundation model on the diffusion transformer backbone (peebles2023scalable; liu2023vit) with flow matching that takes video inputs and text conditioning to generate audio outputs. It undergoes standard pre-training on large-scale video-audio pairs to establish basic generation capabilities. While this architecture provides a solid foundation for V2A generation, it has two critical limitations that hinder effective multi-dimensional CoT reasoning: insufficient video understanding for complex and diverse scenarios, and limited text processing capabilities for structured reasoning content. Therefore, we enhance the ThinkSound architecture (liu2025thinksound) with the following two modifications to facilitate multi-dimensional CoT reasoning.
VideoPrism for Enhanced Video Understanding. Most existing V2A models, including ThinkSound, adopt CLIP-based image encoders (radford2021learning) that process video frames independently as static images. This approach lacks comprehensive video understanding and fails to handle complex, diverse video scenarios in real-world applications. We replace CLIP with VideoPrism (zhao2024videoprism), a state-of-the-art (SOTA) video encoder pre-trained on large-scale video data. VideoPrism employs a unified vision transformer architecture specially designed for video understanding, capturing rich semantic representations of objects, actions, and environmental contexts that are crucial for our multi-dimensional reasoning modules.
T5-Gemma for CoT-Aware Text Encoding. Our CoT modules produce analytical text containing logical structures, causal relationships, and multi-faceted reasoning patterns that require sophisticated language understanding; yet the standard T5 encoders in ThinkSound struggle with the complex, structured reasoning text generated by our CoT modules. Hence, we make another essential enhancement by upgrading the T5 encoder to T5-Gemma (zhang2025encoder). T5-Gemma adapts the reasoning capabilities of decoder-only LLMs into an efficient encoder-decoder architecture, effectively enabling proper conditioning on our structured CoTs for the generation model with its stronger reasoning comprehension capabilities.
@@ 3.2 Decomposing Multi-Dimensional CoT Reasoning
To achieve superior reasoning capabilities simultaneously across all four dimensions, we first employ Gemini 2.5 Pro (comanici2025gemini) for CoT data construction, leveraging its outstanding multimodal understanding and strong reasoning capabilities. Next, using this high-quality training data, we fine-tune the highly competitive open-source video language model VideoLLaMA2 (cheng2024videollama) to generate four specialized CoTs: Semantic CoT identifies audio events and their characteristics from audio-visual content; Temporal CoT determines the sequential ordering of audio events; Aesthetic CoT focuses on audio quality aspects like naturalness and fidelity; and Spatial CoT analyzes sound positioning, including directional placement and distance for proper spatialization. The four specialized CoTs are then concatenated in this order to form the multi-dimensional CoT (as depicted in Figure 1) and used as enhanced, structured text conditioning to fine-tune our audio foundation model (Section 3.1), enabling the model to learn from explicit reasoning patterns and acquire better generalization by understanding the underlying logic behind audio-visual correspondences.
@@ 3.3.1 Multi-dimensional Reward Functions
To address this limitation, we design four specialized reward functions that align with our CoT dimensions: Semantic Reward, measured by MS-CLAP (elizalde2024natural), a commonly used audio-text alignment model for evaluating content similarity; Temporal Reward, assessed via Synchformer (iashin2024synchformer), a highly competitive model specifically designed to detect audio-visual synchrony; Aesthetic Reward, which uses the leading-edge assessor for audio aesthetic quality, Meta Audiobox Aesthetics (tjandra2025meta), as a no-reference model trained to predict human Mean Opinion Scores (MOS); and Spatial Reward, which employs the high-performing StereoCRW (chen2022sound) to verify directional positioning accuracy. This multi-objective approach allows for a balanced and comprehensive optimization across all key perceptual dimensions.
@@ 3.3.2 Fast-GRPO with Random-Window Exploration
To align the audio foundation model with the multi-dimensional human preference, we adopt GRPO for its stability. While generation of our flow matching model is inherently deterministic (an ODE), it can be equivalently formulated as a stochastic process (an SDE) that enables RL-based optimization (see Appendix B.1 for details). Prior works (xue2025dancegrpo; liu2025flow) construct the Markov Decision Process (MDP) within the GRPO training process by applying this SDE formulation across the entire denoising trajectory. This “pure SDE” approach, however, forces GRPO to evaluate the policy at every step, creating a significant efficiency bottleneck.
@@
where $A^{i}$ is the group-normalized advantage for the $i$-th sample (Eq. 6). This design reduces the policy-model NFE (Number of Function Evaluations) from $T$ to $w$ per sample, yielding a near-linear complexity of GRPO training. We notice that some contemporaneous works, such as MixGRPO (li2025mixgrpo), also propose hybrid ODE-SDE. Considering the concurrency of research and their differences from Fast-GRPO on window design, modalities, and scopes, our Fast-GRPO is a valid innovation for enabling efficient multi-dimensional RL training of diffusion models.
@@ 4.1 Experimental Setup
Evaluation Metrics. We conduct comprehensive evaluations using both objective and subjective metrics to assess the four key perceptual dimensions. For objective evaluation, we adopt established metrics across multiple dimensions. Following ThinkSound, we employ CLAP score for text-audio semantic alignment, DeSync measured by Synchformer for video-audio temporal synchrony, Fréchet Distance (FD(kilgour2018fr) in the VGGish feature space, and Kullback-Leibler (KL) Divergence (copet2024simple) based on predictions from the PaSST model for audio distribution similarity liu2025thinksound. For spatial accuracy of generated stereo audio, we adopt GCC MSE and CRW MSE sun2024both to evaluate both difference of arrival (DoA) and interaural time difference (ITD). To measure aesthetic quality, we evaluate production quality (PQ), production complexity (PC), content enjoyment (CE), and content usefulness (CU) scores from Audiobox-Aesthetics (tjandra2025meta). For subjective evaluation, we employ Mean Opinion Score (MOS) across two complementary dimensions: MOS-Q (Quality) evaluates the aesthetic quality and audio fidelity of generated audio, while MOS-C (Consistency) evaluates the comprehensive alignment between audio and video, encompassing semantic consistency, temporal synchrony, and spatial accuracy. More details of evaluation metrics are in Appendix F.
Implementation Details are in Appendix D. Since the multi-dimensional CoT fine-tuning and the RL post-training are based on VGGSound, evaluations on the VGGSound test set are in-domain evaluations. Competitive baselines include Frieren (16k, mono) (wang2024frieren), V2A-Mapper (16k, mono) (wang2024v2a), AudioX (44k, stereo) (tian2025audiox), HunyuanVideo-Foley (44k, mono) (shan2025hunyuanvideo), MMAudio (44k, mono) (cheng2024taming), and ThinkSound (44k, stereo) (liu2025thinksound).
@@ Appendix A Related Work on Video-to-Audio Generation
V2A generation has recently gained significant attention with advances in multimodal AI systems (simeoni2025dinov3; brooks2024video; team2024gemini). Current V2A methods predominantly employ latent diffusion models (xing2024seeing; liu2025omniaudio; wang2024frieren), while some explore autoregressive token-based approaches. Diff-Foley (luo2023diff) and VTA-LDM (xu2024video) represent standard latent diffusion models conditioned on video features, while FoleyGen (mei2024foleygen) and V-AURA (viertola2025temporally) frame conditional audio generation as next-token prediction using visual features. Early methods focus on improving semantic consistency through better video representations. Works like CAVP (luo2023diff) and CLIP4CLIP (luo2022clip4clip) employ contrastive learning for video encoding, while adapter-based approaches like FoleyCrafter (zhang2024foleycrafter) build upon pre-trained text-to-audio models (liu2023audioldm; liu2024flashaudio; liu2024audiolcm) for enhanced controllability. Recent advances introduce explicit multimodal conditioning to address semantic limitations. MovieGen Audio (polyak2024movie) conditions on both text and video to generate video-aligned audio, achieving substantial progress and inspiring subsequent video-text-audio generation works (cheng2024taming; chen2025video; shan2025hunyuanvideo; tian2025audiox). Most recently, ThinkSound (liu2025thinksound) innovatively introduces CoT reasoning via MLLMs, replacing simple text prompts with structured reasoning that significantly improves interpretability and narrative coherence. However, existing V2A methods suffer from objective entanglement—optimizing competing perceptual goals through a single reconstruction loss—and lack human preference alignment beyond textual matching. In contrast, we propose PrismAudio, the first reinforcement learning framework for V2A generation with specialized multi-dimensional CoT-reward correspondence to address these fundamental limitations.